# FERGI: Automatic Annotation of User Preferences for Text-to-Image Generation from Spontaneous Facial Expression Reaction

### v2

@@ Abstract
Researchers have proposed to use data of human preference feedback to fine-tune text-to-image generative models. However, the scalability of human feedback collection has been limited by its reliance on manual annotation. Therefore, we develop and test a method to automatically annotate user preferences from their spontaneous facial expression reaction to the generated images. We collect a dataset of Facial Expression Reaction to Generated Images (FERGI) and show that the activations of multiple facial action units (AUs) are highly correlated with user evaluations of the generated images. Specifically, AU4 (brow lowerer) is reflective of negative evaluations of the generated image whereas AU12 (lip corner puller) is reflective of positive evaluations. These can be useful in two ways. Firstly, we can automatically annotate user preferences between image pairs with substantial difference in these AU responses with an accuracy significantly outperforming state-of-the-art scoring models. Secondly, directly integrating the AU responses with the scoring models improves their consistency with human preferences. Finally, this method of automatic annotation with facial expression analysis can be potentially generalized to other generation tasks. The code is available at https://github.com/ShuangquanFeng/FERGI, and the dataset is also available at the same link for research purposes.
@@ 1 Introduction
However, the scalability of human feedback collection has still been limited by its reliance on manual annotation. Therefore, we develop and test a system to automatically annotate user preferences for text-to-image generation from their spontaneous facial expression reaction to the generated images. This automatic annotation requires zero additional effort from the real users of text-to-image generators and is thus expected to increase the scalability of human preference feedback data collection.
We present the Facial Expression Reaction to Generated Images (FERGI) dataset, which comprises video recordings of 33 participants’ facial expression reaction to 2827 images generated by Stable Diffusion (SD) v1.4 [58] based on 576 different self-drafted text prompts, along with their feedback on the generated images from manual input. We estimate the activation of facial action units (AUs), as defined in the Facial Action Coding System (FACS, a comprehensive system breaking down facial expressions into individual components of AUs of muscle movements [20]), in the facial expression reaction videos with a pre-trained model and show that the activations of multiple AUs are highly correlated with user evaluations of the generated images. We further propose a method of automatically annotating user preferences by detecting the activations of AU4 (brow lowerer, see Fig. 5) and AU12 (lip corner puller, see Fig. 6) in the facial expression reaction to the generated images.
Figure 1: A flow chart summarizing our proposed pipeline. The solid lines represent the procedures demonstrated in the paper while the dashed lines represent the procedures proposed but not demonstrated. Specifically, we developed a model and method for computing an AUcomb valence score and integrating it with state-of-the-art text-to-image scoring models. We propose that the AUcomb valence score could be used to fine-tune/improve these text-to-image scoring models and that the integrated ensemble score could be used to fine-tune the text-to-image generative models.
To better illustrate the idea of our paper, we present Fig. 1 showing a flow chart summarizing the pipeline the paper proposed.
@@ 2 Related Work
Facial Expression Recognition and Applications. Automatic facial expression recognition has advanced rapidly in recent years [42]. As emotional responses to image generation can be more complicated than single categories of the most widely-researched seven basic emotions [19, 18, 47], we decided to analyze the facial expression reactions directly in terms of muscle movements, as defined by AUs in FACS [20]. AU detection and estimation has attracted increasing interest [27, 61, 84, 83, 11, 43, 44, 69, 34, 35, 85, 46, 65, 66] and been used for facial emotion recognition [77, 79, 78, 76]. Specifically related to our usage of AU4 and AU12 to infer user evaluations of generated images, researchers have shown that the activation of the corrugator (activated in AU4) and zygomaticus (activated in AU12) are associated with negative and positive emotions respectively and are the most investigated muscle activations in studying affect and emotion [72]. Also, the activity of the corrugator supercilii (activated in AU4) is positively associated with amygdala and negatively associated with ventromedial prefrontal cortex activity [24] and increases with negative stimuli, such as negative images [6, 5, 38, 37]. For applications, [3, 71] showed the effectiveness of using facial expression to infer/analyze human preferences.
@@ 3.1 Participants
39 participants were recruited from the SONA system of University of California San Diego (UCSD) and completed the study asynchronously on personal computers. 6 participants were removed from the dataset for various reasons, including failure to participate, failure in video recording, and not permitting sharing of their video recordings for research. Therefore, we ended up with a dataset of 33 participants that is available upon request for research purposes at https://github.com/ShuangquanFeng/FERGI. The collection and usage of human participant data in the dataset were approved by the Institutional Review Board (IRB) of UCSD.
@@ 3.2.1 Prompt Creation
There are two types of sessions, structured input and free-form input. In both types, the participant can freely draft any text prompt as the input as long as it is not NSFW. The only difference between the two types lies in how the input text prompt is created. In a structured input session, the participant fills out a form to separately specify different elements desired in the images to generate: animate objects, inanimate objects, interactional relations (between two objects), positional relations (between two objects), location, style (of the image), and keywords. Based on the contents of the form, a text prompt is then automatically generated by the Large Language Model (LLM) OpenAI gpt3.5-turbo. The participant can freely adjust the generated prompt as the final input text prompt. In a free-form input, the participant directly enters the entire final input text prompt.
@@ 3.2.3 Image Presentation, Annotation, and Ranking
After the annotation of the last image, the participant ranks the 5 generated images from best to worst, and the data of the session is uploaded to the web server. That marks the end of the session.
@@ 4 AU Model Training
We analyze the facial expression reaction of the users by estimating the activation of their AUs. We trained our AU estimation model on the DISFA [50, 49] and DISFA+ [48] datasets. They are among the largest, most popular, and most cited AU datasets with intensity annotations. In these datasets, each frame is manually annotated by a human expert with intensities on a scale of 0 to 5, for activation of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop). Video frames are first preprocessed with face detection [45], face alignment [7], and a combination of histogram equalization and linear mapping [36] before being fed into the model for training, testing, or inference. For model training, we used the neural network IR-50 [12] pre-trained on Glink360k [1] and fine-tuned it on the DISFA and DISFA+ datasets [50, 48]. A single network learned the estimation of all labeled AUs with both regression and ordinal classification [52]. More details of the training of the AU model is elaborated in the Supplementary Materials.
@@ 5 Facial Feature Extraction
In this section, we elaborate the procedure of extracting facial features from the video clips associated with each generated image based on the trained AU recognition model. Similar procedures were used to extract features from baseline clips and reaction clips. Further details are in the Supplemental Materials.
@@ 5.1 Data Filtering
Robustness against face occlusion and faces with non-frontal poses is an ongoing challenge for facial expression recognition models [81, 28]. Our AU model was trained on datasets consisting predominantly of frontal facial images without occlusion and thus for each 5-second video clip, we exclude frames with a low face detection confidence score (given by MediaPipe [45]) and frames with an off-angle pitch or yaw (estimated based on detected facial landmarks [7]). Then, if a clip has more than $20\%$ of its total frames being excluded, the whole clip, and its associated data (survey and ranking results) are excluded from further analysis.
@@ 5.2 AU Activation Value
We compute an AU activation value $\alpha_{i}$ for each trained AUi (with $i$ representing the index of the AU) in each video clip. The AU activation value is computed as follows:
  • Firstly, we estimate the intensity of each AU for each frame of the video clip with the AU model: $\hat{y}_{i,\mathrm{reg}}^{(1)},\hat{y}_{i,\mathrm{reg}}^{(2)},\dots$, where the superscript indicates the index of the frame.
  • Secondly, we compute the moving window mean of the estimated AU intensities for every 0.1 seconds (3 frames for 30 FPS):
@@
  • Finally, we compute the AU activation of the clip as the difference between the maximum AU intensity within the video clip and the AU intensity of the first 0.1 seconds:
@@ 6.1 Statistical Analysis
(a) disappointed
(b) satisfied
(c) surprised
(d) disgusted
(e) amused
(f) scared
Figure 4: Reported emotions highly correlated with AU activation values. The distributions of the activation values of multiple AUs for images eliciting different emotions of the participants as reported in answers to the question “Did you feel any of the following emotions when you saw the image?” Each subfigure shows the results for a different emotion (indicated in the subcaptions) and contains multiple pairs of boxplots representing the results for different AUs. The x-axis represents the indices of the AUs while the y-axis represents the AU activation values in the reaction clip of each image. The parenthesized numbers in the legends represent the number of responses of that type. In each pair of boxplots, the boxplot on the left/right side represents the AU activation values for images that did/didn’t elicit the corresponding emotion. In each boxplot, the bottom of the box represents the first quartile (25th percentile) of the AU activation values, the top of the box represents the third quartile (75th percentile), and the line in the middle of the box represents the median.
@@
  • Emotions. For the question “Did you feel any of the following emotions when you saw the image?”, the participant can independently select 0 to 6 answers from the six options: disappointed, satisfied, surprised, disgusted, amused, and scared. Each of the reported emotions was found to be significantly correlated with the activation values of multiple AUs.
    • Disappointed. “Disappointing” images are associated with lower activation values of AU1, AU2, and AU12 and higher activation values of AU4 as shown in Fig. 3(a) (Wilcoxon rank-sum tests: $z\approx-4.25$ and $p<5\times 10^{-5}$ for AU1, $z\approx-5.82$ and $p<1\times 10^{-8}$ for AU2, $z\approx 4.82$ and $p<5\times 10^{-6}$ for AU4, $z\approx-4.86$ and $p<5\times 10^{-6}$ for AU12).
    • Satisfied. “Satisfying” images are associated with lower activation values of AU4, AU5, AU9, and AU25 as shown in Fig. 3(b) (Wilcoxon rank-sum tests: $z\approx-9.07$ and $p<5\times 10^{-19}$ for AU4, $z\approx-6.62$ and $p<5\times 10^{-11}$ for AU5, $z\approx-6.93$ and $p<5\times 10^{-12}$ for AU9, $z\approx-3.79$ and $p<5\times 10^{-4}$ for AU25).
@@
    • Surprised. “Surprising” images are associated with higher activation values of AU9, AU12, AU25, and AU26 as shown in Fig. 3(c) (Wilcoxon rank-sum tests: $z\approx 3.78$ and $p<5\times 10^{-4}$ for AU9, $z\approx 4.29$ and $p<5\times 10^{-5}$ for AU12, $z\approx 4.01$ and $p<1\times 10^{-4}$ for AU25, $z\approx 4.45$ and $p<1\times 10^{-5}$ for AU26).
@@
    • Disgusted. “Disgusting” images are associated with higher activation values of AU4, AU5, and AU9 as shown in Fig. 3(d) (Wilcoxon rank-sum tests: $z\approx 8.43$ and $p<1\times 10^{-16}$ for AU4, $z\approx 3.94$ and $p<1\times 10^{-4}$ for AU5, $z\approx 4.87$ and $p<1\times 5^{-6}$ for AU9).
@@
    • Amused. “Amusing” images are associated with higher activation values of AU6, AU9, AU12, AU20, and AU25 as shown in Fig. 3(e) (Wilcoxon rank-sum tests: $z\approx 7.00$ and $p<1\times 10^{-11}$ for AU6, $z\approx 5.01$ and $p<1\times 10^{-6}$ for AU9, $z\approx 9.15$ and $p<1\times 10^{-19}$ for AU12, $z\approx 6.12$ and $p<1\times 10^{-9}$ for AU20, $z\approx 3.85$ and $p<1\times 10^{-4}$ for AU25).
@@
    • Scared. “Scary” images are associated with higher activation values of AU4 and AU5 as shown in Fig. 3(f) (Wilcoxon rank-sum tests: $z\approx 5.60$ and $p<5\times 10^{-8}$ for AU4, $z\approx 4.39$ and $p<5\times 10^{-5}$ for AU5).
@@
Note that the results presented above selectively include the AUs whose associated p-values (as reported before corrections) are below the Bonferroni-corrected significance threshold of $0.05/[12\times(2+6)]$, with 12 representing the 12 investigated AUs, 2 representing the two tests for overall ratings (1 for the raw rating and 1 for the extremity of the rating), and 6 representing the 6 tests for 6 emotions.
6.2 Automatic Annotation of User Preferences between Image Pairs Using AU4 and AU12
Figure 5: An example of AU4 activation in response to a low-quality image generation. (a) The input text prompt. (b) The facial expression of the participant before seeing the generated image. (c) The generated image. (d) The facial expression of the participant after seeing the generated image (the frame with the highest estimated AU4 intensity). In the annotation survey, the participant gave an overall rating of 2, an image-text alignment rating of 2, and a fidelity rating of 1; she identified the image with the issues “output contains unwanted content that was not mentioned in the text prompt” and “existence of body problem”; she reported that she felt “disappointed” when seeing the image; she ranked the image as the worst among the five images generated from the same prompt.
Figure 6: An example of AU12 activation in response to a high-quality image generation. (a) The input text prompt. (b) The facial expression of the participant before seeing the generated image. (c) The generated image. (d) The facial expression of the participant after seeing the generated image (the frame with the highest estimated AU12 intensity). In the annotation survey, the participant gave ratings of 7 for each of image-text alignment, fidelity, and overall rating; he did not identify any issues with the image; he reported feeling “satisfied”, “surprised”, and “amused” when seeing the image; he ranked the image as the best among the five images generated from the same prompt.
We aim to use the estimated AU activation values to automatically annotate user preferences between images generated from the same input text prompt. We have observed that the activations of AU2, AU4, AU9, and AU12 are significantly correlated with the the raw overall ratings in the previous subsection. After comparing combinations between them, we developed the following pipeline of using AU4 and AU12 to do automatic annotation (examples of AU4 and AU12 responses are shown in Figs. 5 and 6). (Inclusion of AU2 and AU9 did not help.) We also did not find a way to effectively incorporate the information conveyed by AUs that correlate only with the extremity of the ratings. Collection of a much larger-scale dataset in the future might enable effective incorporation of such information with a more complex model.
6.2.1 Participant Exclusion
While estimations of AU12 are generally reliable and stable, multiple factors may result in unreliable AU4 estimation from a small number of people by our AU model. We excluded 3 participants with unreliable AU4 estimation due to bangs covering the area between the eyebrows (2 participants) or highly unstable camera angles (1 participant).
6.2.2 AUcomb Valence Score
The AU4 activation value $\alpha_{4}$ and AU12 activation value $\alpha_{12}$ are good indicators of the user’s evaluation of the generated image, but the raw values might not linearly reflect the associated evaluation. Therefore, we propose an AU4 valence score $s_{\mathrm{AU4}}$ and an AU12 valence score $s_{\mathrm{AU12}}$, designed to have a more linear relationship and positive correlation with the associated evaluation of the generated image. We define the $s_{\mathrm{AU4}}$ as the negative of an exponential decay function of AU4 activation value $\alpha_{4}$ and $s_{\mathrm{AU12}}$ as the positive exponential decay function of AU12 activation value $\alpha_{12}$ as follows:
$s_{\mathrm{AU4}}=-(1-e^{-k\cdot\alpha_{4}})$ (3)
and
$s_{\mathrm{AU12}}=(1-e^{-k\cdot\alpha_{12}}),$ (4)
where $k$ is an exponential decay coefficient to fit. We further define the final AUcomb valence score $s_{\mathrm{AUcomb}}$ as the linear combination of them:
$s_{\mathrm{AUcomb}}=w_{\mathrm{AU4}}s_{\mathrm{AU4}}+w_{\mathrm{AU12}}s_{\mathrm{AU12}}$ (5)
subject $\text{subject to}\quad w_{\mathrm{AU4}}+w_{\mathrm{AU12}}$ $=1,$ (6)
where $w_{\mathrm{AU4}}$ and $w_{\mathrm{AU12}}$ represent the weights for AU4 and AU12 respectively.
6.2.3 Baseline Scores
To compare the performance of our AU valence score, we use the CLIP score [55], Aesthetic score [60], BLIP score [41], ImageReward score [75], PickScore [33], and HPS v2 score [73] as baselines.

### v3

# FERGI: Automatic Scoring of User Preferences for Text-to-Image Generation from Spontaneous Facial Expression Reaction

@@ Abstract
Researchers have proposed to use data of human preference feedback to fine-tune text-to-image generative models. However, the scalability of human feedback collection has been limited by its reliance on manual annotation. Therefore, we develop and test a method to automatically score user preferences from their spontaneous facial expression reaction to the generated images. We collect a dataset of Facial Expression Reaction to Generated Images (FERGI) and show that the activations of multiple facial action units (AUs) are highly correlated with user evaluations of the generated images. We develop an FAU-Net (Facial Action Units Neural Network), which receives inputs from an AU estimation model, to automatically score user preferences for text-to-image generation based on their facial expression reactions, which is complementary to the pre-trained scoring models based on the input text prompts and generated images. Integrating our FAU-Net valence score with the pre-trained scoring models improves their consistency with human preferences. This method of automatic annotation with facial expression analysis can be potentially generalized to other generation tasks. The code is available at https://github.com/ShuangquanFeng/FERGI, and the dataset is also available at the same link for research purposes.
@@ 1 Introduction
However, the scalability of human feedback collection has still been limited by its reliance on manual annotation. Researchers have used the CLIP score [59], Aesthetic score [64], BLIP score [43], ImageReward score [79], PickScore [35], and HPS v2 score [77] to automatically score human preferences of text-to-image generation based on the input text prompts and the generated images. In this study, we develop and test a system to automatically score user preferences from their spontaneous facial expression reaction to the generated images, which is complementary to the models based only on the texts and images. This automatic scoring requires zero additional effort from the real users of text-to-image generators.
We present the Facial Expression Reaction to Generated Images (FERGI) dataset, which comprises video recordings of 33 participants’ facial expression reaction to 2827 images generated by Stable Diffusion (SD) v1.4 [62] based on 576 different self-drafted text prompts, along with their feedback on the generated images from manual input. We estimate the activation of facial action units (AUs), as defined in the Facial Action Coding System (FACS, a comprehensive system breaking down facial expressions into individual components of AUs of muscle movements [21]), in the facial expression reaction videos with a model we trained on external datasets and show that the activations of multiple AUs are highly correlated with user evaluations of the generated images. We further propose a method of automatically scoring user preferences by detecting the activations of the AUs in the facial expression reaction to the generated images and feeding them into a neural network.
@@ 2 Related Work
Facial Expression Recognition and Applications. Automatic facial expression recognition has advanced rapidly in recent years [44]. As emotional responses to image generation can be more complicated than single categories of the most widely-researched seven basic emotions [20, 19, 50], we decided to analyze the facial expression reactions directly in terms of muscle movements, as defined by AUs in FACS [21]. AU detection and estimation has attracted increasing interest [29, 65, 88, 87, 12, 45, 46, 73, 36, 37, 89, 49, 69, 70] and been used for facial emotion recognition [81, 83, 82, 80]. For example, researchers have shown that the activation of the corrugator (activated in AU4) and zygomaticus (activated in AU12) are associated with negative and positive emotions respectively and are the most investigated muscle activations in studying affect and emotion [76]. Also, the activity of the corrugator supercilii (activated in AU4) is positively associated with amygdala and negatively associated with ventromedial prefrontal cortex activity [26] and increases with negative stimuli, such as negative images [7, 6, 40, 39]. For applications, [4, 75] showed the effectiveness of using facial expression to infer/analyze human preferences.
@@ 3.1 Participants
39 participants were recruited from the SONA system of University of California San Diego (UCSD) and completed the study asynchronously on personal computers. 6 participants were removed from the dataset for various reasons, including failure to participate, failure in video recording, and not permitting sharing of their video recordings for research. This resulted in a dataset of 33 participants that will be made available for research. The collection and usage of human participant data in the dataset were approved by the Institutional Review Board (IRB) of UCSD.
@@ 3.2.1 Prompt Creation
There are two types of sessions, structured input and free-form input. In both types, the participant can freely draft any text prompt as long as it is not NSFW. The only difference between the two types lies in how the input text prompt is created. In a structured input session, the participant fills out a form to separately specify different elements desired in the images to generate: animate objects, inanimate objects, interactional relations (between two objects), positional relations (between two objects), location, style (of the image), and keywords. Based on the contents of the form, a text prompt is then automatically generated by the Large Language Model (LLM) OpenAI gpt3.5-turbo. The participant can freely adjust the generated prompt as the final input text prompt. In a free-form input, the participant directly enters the entire final input text prompt.
@@ 3.2.3 Image Presentation, Annotation, and Ranking
After the annotation of the last image, the participant ranks the 5 generated images from best to worst, and the data of the session is uploaded to the web server ending the session.
@@ 4 AU Model Training
We analyze the facial expression reaction of the users by estimating the activation of their AUs. We trained our AU estimation model on the DISFA [53, 52] and DISFA+ [51] datasets. They are among the largest, most popular, and most cited AU datasets with intensity annotations. In these datasets, each frame is manually annotated by a human expert with intensities on a scale of 0 to 5, for activation of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop). (See Fig. S10 for a visual reference guide for the analyzed AUs.) Video frames are first preprocessed with face detection [48], face alignment [8], and a combination of histogram equalization and linear mapping [38] before being fed into the model for training, testing, or inference. For model training, we used the neural network IR-50 [13] pre-trained on Glink360k [1] and fine-tuned it on the DISFA and DISFA+ datasets [53, 51]. A single network learned the estimation of all labeled AUs with both regression and ordinal classification [55]. More details of the training of the AU model is elaborated in the Appendix / Supplementary Materials.
The comparison between our AU estimation model’s performance on the DISFA dataset and other state-of-the-art models, including OpenFace 2.0 [2], CCNN-IT [73], 2DC [47], SCC-Heatmap [23], and iARL [65] is shown in Tab. 1. Our model’s Pearson’s $r$ (higher is better) is close to that of OpenFace 2.0; our model’s ICC(3,1) (higher is better) is the best among all models. Although our model’s MAE (lower is better) is less competitive among all models, we believe that ICC(3,1) is a better metric for evaluating the performance of AU estimation models on the DISFA dataset considering the high imbalance of the dataset.
Table 1: Comparison of our AU estimation model’s performance on the DISFA dataset with other models
AU 1 2 4 5 6 9 12 15 17 20 25 26 Avg.
Pearson’s $r$ (higher is better)
OpenFace 2.0 .64 .50 .70 .67 .59 .54 .85 .39 .49 .22 .85 .67 .59
Our Model .53 .46 .74 .57 .56 .60 .85 .42 .41 .21 .94 .66 .58
ICC(3,1) (higher is better)
CCNN-IT .18 .15 .61 .07 .65 .55 .82 .44 .37 .28 .77 .54 .45
2DC .70 .55 .69 .05 .59 .57 .88 .32 .10 .08 .90 .50 .50
SCC-Heatmap .73 .44 .74 .06 .27 .51 .71 .04 .37 .04 .94 .78 .47
iARL .13 .36 .68 .22 .56 .36 .86 .52 .37 .12 .96 .60 .48
Our Model .53 .46 .74 .56 .55 .57 .85 .41 .40 .20 .93 .66 .57
MAE (lower is better)
CCNN-IT 87 .63 .86 .26 .73 .57 .55 .38 .57 .45 .81 .64 .61
SCC-Heatmap .16 .16 .27 .03 .25 .13 .32 .15 .20 .09 .30 .32 .20
iARL .30 .31 .52 .04 .36 .30 .31 .05 .33 .08 .29 .26 .26
Our Model .37 .39 .50 .14 .37 .21 .33 .24 .47 .23 .32 .42 .33
@@ 5 Facial Feature Extraction
In this section, we elaborate the procedure of extracting facial features from the video clips associated with each generated image based on the trained AU recognition model. Similar procedures were used to extract features from baseline clips and reaction clips. Further details are in the Appendix / Supplementary Materials.
@@ 5.1 Data Filtering
Robustness against face occlusion and faces with non-frontal poses is an ongoing challenge for facial expression recognition models [85, 30]. Our AU model was trained on datasets consisting predominantly of frontal facial images without occlusion and thus for each 5-second video clip, we exclude frames with a low face detection confidence score (given by MediaPipe [48]) and frames with an off-angle pitch or yaw (estimated based on detected facial landmarks [8]). If a clip has more than $20\%$ of its total frames excluded, the whole clip, and its associated data (survey and ranking results) are excluded from further analysis.
@@ 5.2 AU Activation Value
We compute an AU activation value $\alpha_{i}$ for each trained AUi in each video clip. The AU activation value is computed as follows:
  • Firstly, the intensity of each AU for each frame of the video clip is estimated with the AU model: $\hat{y}_{i,\mathrm{reg}}^{(1)},\hat{y}_{i,\mathrm{reg}}^{(2)},\dots$, where the superscript indicates the index of the frame.
  • Secondly, a moving window mean of the estimated AU intensities for every 0.1 seconds (3 frames for 30 FPS) is computed:
@@
  • Finally, the AU activation of the clip as the difference between the maximum AU intensity within the video clip and the AU intensity of the first 0.1 seconds is computed:
@@ 6.1 Statistical Analysis
(a) surprised
(b) disgusted
(c) amused
(d) scared
(e) disappointed
(f) satisfied
Figure 3: Reported emotions highly correlated with AU activation values. The distributions of the activation values of multiple AUs for images eliciting different emotions of the participants as reported in answers to the question “Did you feel any of the following emotions when you saw the image?”. Each subfigure shows results for a different emotion (indicated in the captions) and contains multiple pairs of boxplots representing the results for different AUs. The indices of the AUs are along the x-axis, while the y-axis represents the AU activation values in the reaction clip of each image. The parenthesized numbers in the legends give the number of responses of that type. In each pair of boxplots, the boxplot on the left/right side represents the AU activation values for images that did/didn’t elicit the corresponding emotion. In each boxplot, the bottom/top of the box represents the first/third quartile (25th/75th percentile) of the AU activation values, and the line in the middle of the box represents the median.
@@
  • Emotions. For the question “Did you feel any of the following emotions when you saw the image?”, the participant can independently select 0 to 6 answers from the six options: surprised, disgusted, amused, scared, disappointed, and satisfied. Each of the reported emotions was found to be significantly correlated with the activation values of multiple AUs.
@@
    • Surprised. “Surprising” images are associated with higher activation values of AU9, AU12, AU25, and AU26 as shown in Fig. 3(a) (Wilcoxon rank-sum tests: $z\approx 3.78$ and $p<5\times 10^{-4}$ for AU9, $z\approx 4.29$ and $p<5\times 10^{-5}$ for AU12, $z\approx 4.01$ and $p<1\times 10^{-4}$ for AU25, $z\approx 4.45$ and $p<1\times 10^{-5}$ for AU26).
@@
    • Disgusted. “Disgusting” images are associated with higher activation values of AU4, AU5, and AU9 as shown in Fig. 3(b) (Wilcoxon rank-sum tests: $z\approx 8.43$ and $p<1\times 10^{-16}$ for AU4, $z\approx 3.94$ and $p<1\times 10^{-4}$ for AU5, $z\approx 4.87$ and $p<1\times 5^{-6}$ for AU9).
@@
    • Amused. “Amusing” images are associated with higher activation values of AU6, AU9, AU12, AU20, and AU25 as shown in Fig. 3(c) (Wilcoxon rank-sum tests: $z\approx 7.00$ and $p<1\times 10^{-11}$ for AU6, $z\approx 5.01$ and $p<1\times 10^{-6}$ for AU9, $z\approx 9.15$ and $p<1\times 10^{-19}$ for AU12, $z\approx 6.12$ and $p<1\times 10^{-9}$ for AU20, $z\approx 3.85$ and $p<1\times 10^{-4}$ for AU25).
@@
    • Scared. “Scary” images are associated with higher activation values of AU4 and AU5 as shown in Fig. 3(d) (Wilcoxon rank-sum tests: $z\approx 5.60$ and $p<5\times 10^{-8}$ for AU4, $z\approx 4.39$ and $p<5\times 10^{-5}$ for AU5).
    • Disappointed. “Disappointing” images are associated with lower activation values of AU1, AU2, and AU12 and higher activation values of AU4 as shown in Fig. 3(e) (Wilcoxon rank-sum tests: $z\approx-4.25$ and $p<5\times 10^{-5}$ for AU1, $z\approx-5.82$ and $p<1\times 10^{-8}$ for AU2, $z\approx 4.82$ and $p<5\times 10^{-6}$ for AU4, $z\approx-4.86$ and $p<5\times 10^{-6}$ for AU12).
    • Satisfied. “Satisfying” images are associated with lower activation values of AU4, AU5, AU9, and AU25 as shown in Fig. 3(f) (Wilcoxon rank-sum tests: $z\approx-9.07$ and $p<5\times 10^{-19}$ for AU4, $z\approx-6.62$ and $p<5\times 10^{-11}$ for AU5, $z\approx-6.93$ and $p<5\times 10^{-12}$ for AU9, $z\approx-3.79$ and $p<5\times 10^{-4}$ for AU25).
@@
Note that the results presented above selectively include the AUs whose associated p-values (as reported before corrections) are below the Bonferroni-corrected significance threshold of $0.05/[12\times(2+6)]$ (12 investigated AUs, 2 tests for overall ratings (1 for the raw rating and 1 for the extremity of the rating), and 6 emotions).
6.2 Automatic Annotation of User Preferences between Image Pairs
6.2.1 Pre-trained Scoring Models
Researchers have used the CLIP score [59], Aesthetic score [64], BLIP score [43], ImageReward score [79], PickScore [35], and HPS v2 score [77] to estimate human preferences of text-to-image generation based on the input text prompts and the generated images.